Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • 20 Hz implemented and running as of Thursday, 13 May.
  • Wide band tests; plan for RSRO-specific tests this week.
  • Thursday data review meeting; only intermittently hitting level of analysis depth. Provide more advanced warning/schedule to facilitate; proposed draft schedule:
    • 20 May: Rupen: Wideband results/RSRO readiness; Dhawan/Rupen: Reference pointing status
    • 27 May: Sault et al: Anti-aliasing results; Rupen/Perley: Switched power calibration
    • 03 June: Sault/Greisen/Owen: Data weights/noise - REWAY results/status; Wrobel: OST heuristics
    • 10 June - No meeting - summer school - use part of Monday meeting for: Butler: Planetary observing (prep for Cassini); Lazio: Phased array status
    • 17 June - Myers/Moellenbrock Polarization update; Goss/van Gorkom: Bandpass update
    • 24 June - Perley: RFI overview; Rupen/Sowinski/Pokorny: Zeros redux
    • 01 June - Hunter: 3-bit sampler tests; Rupen/Perley: WIDAR acceptance tests
    • 08 July - Hunter/Lyons: Visualizing the Correlator (OPT); Sjouwerman/Greisen/Moellenbrock: Doppler tracking in Post-processing
  • Offer out on EVLA Postdoc position.
  • *Today*: Urvashi's thesis defense at 1500 in the auditorium
  • Staff Travel:
    • Lorant: 9 May - 22 May
    • Steve: 11-18 May
    • Bryan: 17 May - 21 May
    • Claire: 19 - 21 May
    • Joe: 19 - 24 May
    • Emmanuel: 21 May - 7 June
    • Rick: 25-29 May
    • Joe L: 27 May -3 June
    • Michael: 27 May - 1 June
    • Gustaaf: 1 - 4 June
    • Eric: 4 - 19 June

Commissioning review/reports

Notes for week of 10 May 2010 (Black: old but still relevant; Blue: new)

  • Meta summary
    • Multi-config SBs in regular use; still using C-band reference pointing for already verified projects; noted 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
  wide band modes available only in webtest OST (reconciling status database) OPT selected limits not possible
Running 20 Hz
15-Apr CM vers
    - needs check on short pol'n runs Need decision on large (i.e., RSRO) data set distribution media   Operational for OSRO1/OSRO2
30s overhead with configuration changes
CM memory leak requires x'ly restart

Critical Commissioning

System Verification

  • Array/Receivers (Rick)
    • Reference pointing tests planned this week to help quantify status; link from Michael: http://www.vla.nrao.edu/astro/guides/refpt/; status - detailed discussion Thursday.
    • Next try at 20 Hz planned for Tuesday (with/without EA14). Done 13 May.
    • Puzzling numbers in switched power; discussion work this week to understand.
  • WIDAR update (Michael)
    • 5 more CBE nodes to be installed Wednesday (total of 8; moving toward planned design of one BB to each NIC).
    • Sonja plans to test/deploy a new CM which features parallel configuration of BBs; should resolve issues with zeros/overhead of configuration changes. Backed out of; not clear this is the path forward.
    • Issues:
      • Configuration mode changes have overhead of ~30s (Rick); must pad existing SBs to accommodate until resolved; currently holding SBs (rather than editing) in anticipation of this fix. Note: Update from Emmanuel on this - not as dire as we expected - will unblock. https://bugs.aoc.nrao.edu/browse/ECDATA-235
  • Quick Software Systems review:
  • Hardware notes:
    • Scientist reports:
      • ECDATA-234: GvM on AW776: The 13% zeros reported earlier were largely due to zeros in correlation LL on antennas 13, 14, 15, 16. After flagging those the percentage of zeros became 0.3%. Both antennas ea05 and ea25 are absent in scan 17 (which immediately follows a pointing scan). Possibly no valid solution was obtained for these two antennas, but does that flag the subsequent data? And the next scan, which applies the same pointing solutions, has valid data for ea05 and ea25. Bandpass and gain stability are excellent.
      • ECDATA-231: LC on AR708: BP ok; no amp/phase jumps (excellent stability at Ka).
      • ECDATA-229: JMJ on AM991: EA12 IF C/D are weak; EA17 IF B is dead; EA26 IFC no fringes (Fixed); EA17/EA25: Phase difference scatter. EA11 IF2 pol 1 8% amplitude flapping at band edges, 7 degrees in phase.
      • ECDATA-180: LC on AS1039: Q band 41/49 GHz observations loses 41 GHz IF (no fringes); seen after 22 April.
      • ECDATA-227: MK on AS1013: Lots of zeros after reference pointing.
  • OSRO/RSRO/ECSO/TEST/DEMO summary (mostly to be replaced by OST)

Band Checkouts (ECDATA-55) -- No updates.

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

  • P Band Status: %Red%Testing
    • Antennas with no P-band Rx: 7, 16, 22, 24 (no feed); 1, 4, 5 Rxs are not working.
  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated.
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40)
    • L band, HI checkout (ECDATA-47)

  • S Band Status: Ready.
    • S band edge tests (similar result to C for tunings near 4 GHz; potentially some satellite downlink RFI; review of aliasing/out-of-band intermodulation is under review (Rick)
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

  • K Band Status: Ready

  • Q Band Status: Ready.
    • Need Q band suite of tests.
    • Antennas with no Q-band: EA10, EA12, EA17

Observing Mode testing

  • Continuum - ok
  • Spectral Line: (Ott); ECSV-19
    • Bandpass stability: ECSV-35 (on hold due to multiple configuration changes).
    • Bandpass tests: ECSV-92
  • Polarimetry Status: (Myers); ECSV-29
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Combined Low-High Pol checkout test (TPOL0003_sb1378180): Ready for obs (weekend source); pending 30 s overhead issue.
    • Eric reviewing AH1004 data for AIPS
    • Rick: Pol'n tests (receiver rotation) to be scheduled
  • Planetary Bryan away; need planning.
  • Advanced
    • Wideband L (AB1345 - ) Sanjay
    • Solar (Rick) status and path forward. Working with L, Ku (special Rx setups in these two bands). TP tests coming.
    • Phased Array: Joe Lazio will be doing follow-up testing
      • Phased array for pulsars
    • Subarray handling
      • Single-dish VLBI
    • Astrometry
    • OTF Mosaicing

Observing support (e2e capability)

  • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
  • Flux Calibrator Models (Amy); ECSV-43
    • Rupen proposal

Advanced Science Support

OSRO Proposal Testing (ECDATA-34)

RSRO

  • Observations
  • Residents
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests); OSRO data validation
      • This week: -
    • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
      • This week: CASA summer school work
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
      • This week: -
    • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • This week: Phased array tests
    • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • This week: software to store dig spectra; handle SN for Rx
    • Joe Lazio (POC: Claire Chandler): Low frequency observing.
      • This week: phased array tests
    • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • -
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)
      • -
    • ---
    • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).

Actions

  • [Rupen]: Develop daily/post-maintenance day checklists for the system. To be completed as part of stress tests.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at different bands to provide to engineers for review (Q2).
  • [Wrobel/Butler]: Make use of API criteria consistent throughout tools (OPT->OST). List of actions organized. In process.
  • Archive check: Use the SDMset (under Format) to check for any 0 File size BDFs - missing scans/other issues can be trapped this way. Asked John to set this up as an automated check to send an alarm (ACTION).
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting. Butler to report.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

AOB

Planning review (for Q1/Q2 2010):

Note: Beginnings of group pages from the organization page:

-- JosephMcMullin - 17 May 2010
Topic revision: r3 - 2010-05-17, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback